# Constructions, bounds, and algorithms for peaceable queens

### v2

@@ 5. The Algorithm
Figure 5. On the left is an $(a,b)$-plaid $(B,W)$ with $|B|=72$ and $|W|=84$. Let $c_{2}$ and $c_{4}$ be the second and fourth columns, and $r_{24}$ be the bottom row. By swapping on $c_{2}$, then $c_{4}$, and then $r_{24}$, we obtain the peaceful battle $(B^{\prime},W^{\prime})$ on the right. Note that $|B^{\prime}|=|W^{\prime}|=74$, so $\min\{|B^{\prime}|,|W^{\prime}|\}>\min\{|B|,|W|\}$.
Beginning with an initial peaceful battle $(B,W)$, the algorithm performs swaps to ‘improve’ the solution until $\min\{|B|,|W|\}\geqslant q$, where $q$ is an input parameter. There are choices to be made about what ‘improve’ means. One natural candidate is to always improve the parameter $\min\{|B|,|W|\}$ at each step of the algorithm. In practice, we have found that it is advantageous to allow $\min\{|B|,|W|\}$ to decrease, provided $|B|+|W|$ increases. Thus, every swap either increases $\min\{|B|,|W|\}$ or $|B|+|W|$. It appears as if ‘most’ initial choices of $(B,W)$ converge to an optimal solution via a sequence of swaps. In practice, we have observed the fastest convergence by taking $|B|$ to be a random set of size around $\frac{n}{5}$ and $W$ to be $\widehat{B}$. We also allow our algorithm to re-initialize itself using this initial condition. For a precise description of our algorithm, we refer the reader to Algorithm 1.
@@
For the even torus, the algorithm quickly finds an $(a,b)$-plaid, beginning with a peaceful battle $(B,W)$, where $|B|$ only contains the first few cells of the main diagonal. In fact, our algorithm finds solutions which are better than an optimal $(a,b)$-plaid. Therefore, there are optimal solutions for the even torus which are not $(a,b)$-plaids. However, the solutions found by our algorithm have at most two more black queens than an optimal $(a,b)$-plaid. This leads us to the following conjecture.
@@
Finally, for the regular board, our algorithm quickly finds the Ainley construction [1], as well as the other examples in the OEIS which meet the Ainley bound. We did not manage to find any examples that beat the Ainley bound. Thus, our experimental evidence suggests that the Ainley construction is indeed optimal. Although we do not have a proof of correctness, we conjecture that there exist $d\in\mathbb{N}$ and $p>0$ such that our algorithm outputs an optimal peaceful battle $(B,W)$ on $\mathbb{G}_{n}$ after at most $n^{d}$ steps with probability $p$.
@@ Acknowledgements
Thank you to Stephen Moehle for the excellent and thorough code reviews.
@@ Appendix A Best odd torus solutions
(a) $n=13,|B|=16$
(b) $n=15,|B|=20$
(c) $n=17,|B|=28$
(d) $n=19,|B|=32$
(e) $n=21,|B|=40$
(f) $n=23,|B|=48$
(g) $n=25,|B|=56$
(h) $n=27,|B|=66$
(i) $n=29,|B|=76$
(j) $n=31,|B|=88$
(k) $n=33,|B|=101$
(l) $n=35,|B|=110$
(m) $n=37,|B|=126$
(n) $n=39,|B|=144$
(o) $n=41,|B|=156$
(p) $n=43,|B|=162$
(q) $n=45,|B|=184$
(r) $n=47,|B|=197$
(s) $n=49,|B|=212$
(t) $n=51,|B|=252$
(u) $n=53,|B|=250$
(v) $n=55,|B|=266$
(w) $n=57,|B|=285$
(x) $n=59,|B|=304$
(y) $n=61,|B|=324$
(z) $n=63,|B|=348$
Figure 6. Our best odd torus solutions for $13\leqslant n\leqslant 63$.

### v3

@@ 5. The Algorithm
Figure 6. On the left is an $(a,b)$-plaid $(B,W)$ with $|B|=84$ and $|W|=72$. Let $c_{22}$ and $c_{24}$ be the last and third to last columns, and $r_{24}$ be the bottom row. By swapping on $c_{22}$, then $c_{24}$, and then $r_{24}$, we obtain the peaceful battle $(B^{\prime},W^{\prime})$ on the right. Note that $|B^{\prime}|=|W^{\prime}|=74$, so $\min\{|B^{\prime}|,|W^{\prime}|\}>\min\{|B|,|W|\}$.
Beginning with an initial peaceful battle $(B,W)$, the algorithm performs swaps to ‘improve’ the solution until $\min\{|B|,|W|\}\geqslant q$, where $q$ is an input parameter. There are choices to be made about what ‘improve’ means. One natural candidate is to always improve the parameter $\min\{|B|,|W|\}$ at each step of the algorithm. However, in practice we have found that it is advantageous to allow $\min\{|B|,|W|\}$ to decrease, provided $|B|+|W|$ increases. Thus, every swap either increases $\min\{|B|,|W|\}$ or $|B|+|W|$. It appears as if ‘most’ initial choices of $(B,W)$ converge to an optimal solution via a sequence of swaps. In practice, we have observed the fastest convergence by taking $|B|$ to be a random set of size around $\frac{n}{5}$ and $W$ to be $\widehat{B}$. We also allow our algorithm to re-initialize itself using this initial condition. For a precise description of our algorithm, we refer the reader to Algorithm 1.
@@
For the even torus, the algorithm quickly finds an $(a,b)$-plaid. In fact, it is even able to improve on an optimal $(a,b)$-plaid through swapping as seen in Figure 6. Therefore, there are optimal solutions for the even torus which are not $(a,b)$-plaids. However, these ‘swapped’ $(a,b)$-plaids have at most two more black queens than an optimal $(a,b)$-plaid. This leads us to the following conjecture.
@@
Finally, for the regular board, our algorithm quickly finds the Ainley construction [1], as well as other examples (some in the OEIS, some not) which meet the Ainley bound. We did not manage to find any examples that beat the Ainley bound. Thus, our experimental evidence suggests that the Ainley construction is indeed optimal. Although we do not have a proof of correctness, we conjecture that there exist $d\in\mathbb{N}$ and $p>0$ such that our algorithm outputs an optimal peaceful battle $(B,W)$ on $\mathbb{G}_{n}$ after at most $n^{d}$ steps with probability $p$.
@@ Acknowledgements
Matthew Drescher was supported by the National Science Foundation under Grant #2127309 to the Computing Research Association for the CIFellows 2021 Project. Thank you to Stephen Moehle for the excellent and thorough code reviews. We are also grateful to N.J.A. Sloane for reading an early draft of this paper. Matthew Drescher thanks John D. Owens and the Gunrock Graph GPU Lab for introducing him to the problem in a course assignment and for indulging in various interesting discussions. Tony Huynh thanks the Institute for Basic Science in South Korea where part of this research was conducted.
@@ Appendix A Best odd torus solutions
(a) $n=13,\min\{|B|,|W|\}=16$
(b) $n=15,\min\{|B|,|W|\}=20$
(c) $n=17,\min\{|B|,|W|\}=28$
(d) $n=19,\min\{|B|,|W|\}=32$
(e) $n=21,\min\{|B|,|W|\}=40$
(f) $n=23,\min\{|B|,|W|\}=48$
(g) $n=25,\min\{|B|,|W|\}=56$
(h) $n=27,\min\{|B|,|W|\}=66$
(i) $n=29,\min\{|B|,|W|\}=76$
(j) $n=31,\min\{|B|,|W|\}=88$
(k) $n=33,\min\{|B|,|W|\}=101$
(l) $n=35,\min\{|B|,|W|\}=110$
(m) $n=37,\min\{|B|,|W|\}=126$
(n) $n=39,\min\{|B|,|W|\}=144$
(o) $n=41,\min\{|B|,|W|\}=156$
(p) $n=43,\min\{|B|,|W|\}=162$
(q) $n=45,\min\{|B|,|W|\}=184$
(r) $n=47,\min\{|B|,|W|\}=197$
(s) $n=49,\min\{|B|,|W|\}=212$
(t) $n=51,\min\{|B|,|W|\}=252$
(u) $n=53,\min\{|B|,|W|\}=250$
(v) $n=55,\min\{|B|,|W|\}=264$
(w) $n=57,\min\{|B|,|W|\}=285$
(x) $n=59,\min\{|B|,|W|\}=304$
(y) $n=61,\min\{|B|,|W|\}=324$
(z) $n=63,\min\{|B|,|W|\}=348$
(aa) $n=63,\min\{|B|,|W|\}=348$